OpenStudy (anonymous):

how do i solve the equation on the interval 0 le theta < 2pi. cos(2 theta)-3 cos theta + 2 = 0

OpenStudy (anonymous):

anybody can helppp???

OpenStudy (anonymous):

nobody there???

OpenStudy (anonymous):

cos (2 theta) = 2 cos ^2 (theta) -1 replace and treat cos (theta ) as x , solve for x, then, plug back to solve for theta.

OpenStudy (anonymous):

\[\cos 2\theta-3\cos \theta+2=0,2\cos ^{2}\theta-1-3\cos \theta+2=0\] \[2\cos ^{2}\theta-3\cos \theta+1=0\] \[2\cos ^{2}\theta-2\cos \theta-\cos \theta+1=0,\] \[2\cos \theta \left( \cos \theta-1 \right)\] complete it/

OpenStudy (anonymous):

answer has to be one of these: a) \[0, \pi/6, \pi, 11\pi/6\] b) \[\pi/3, \pi/2, 5\pi/3\] c) \[0, \pi/3, 5\pi/3\] d) 0, \[\pi/6, 11\pi/6\]

OpenStudy (anonymous):

Can someone help with this???

OpenStudy (anonymous):

\[2\cos \theta \left( \cos \theta-1 \right)-1\left( \cos \theta-1 \right)=0\] \[\left( \cos \theta-1 \right)\left( 2\cos \theta-1 \right)=0\] \[either \cos \theta-1=0,\cos \theta=1=\cos 0, \in 0\le \theta < 2\pi ,\theta=0 \] \[or 2\cos \theta-1=0,\cos \theta=\frac{ 1 }{2 }=\cos \frac{ \pi }{ 3 },\cos \left( 2\pi-\frac{ \pi }{ 3 } \right)\] \[\theta=\frac{ \pi }{ 3 },\frac{ 5\pi }{ 3 }\]

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

yw